ca you please help find the solution to the following word problem:
Find three consecutive odd integers such that the sum of the first and the third results in 87 less than five times the second

Answer by fractalier(6550)

Let us call the three consecutive odd integers
x, x+2, and x+4
Notice we skip a number because they are odd...
Now translating we get...
x + (x+4) = 5(x+2) - 87
Now solve
2x + 4 = 5x + 10 - 87
2x + 4 = 5x - 77
add 77 to both sides
2x + 81 = 5x
subtract 2x from both sides
81 = 3x
x = 27
x+2 = 29
x+4 = 31